Hello, I have an algebra test on Wednesday and I have no idea of how to go about doing it. Here is a little idea of the type of questions that may be on the test.
Paula has money hidden in a box underneath her dresser. Her friend Jessica wants to know how much is in the "mystery box", so Paula tells her this riddle: The sum of two numbers is 172. The first is eight more than four times the second. If you can find the first number you will know the amount of money in the box.
Do you know how I would go about trying to solve this?

Answer by checkley75(3666)

x+y=172
x=4y+8
(4y+8)+y=172
4y+y=172-8
5y=164
y=164/5 ans.
x+164/5=172
x=172-164/5
x=(5*172-164)/5
x=(860-164)/5
x=696/5 ans.